Find the p -value for the hypothesis test. A random sample of size 51 is taken. The sample has a mean of 392 and a standard deviation of 83. H 0 : µ = 400 H a : µ< 400 The p -value for the hypothesis test is .

Your answer should be rounded to 4 decimal places.
Answer #1

This is the left tailed test .

The null and alternative hypothesis is ,

H0 :  \mu = 400

Ha : \mu <400

Test statistic = z

= (\bar x - \mu ) / \sigma / \sqrt n

= (392 -400) /83 / \sqrt 51

= −0.69

P(z < -0.69 ) = 0.2451

P-value = 0.2451